OpenStudy (anonymous):

Simplify the following expression. a + c − 4b − c + a − b a − 6b a − 5b 2a − 6b 2a − 5b

OpenStudy (anonymous):

combine like terms wanna try it 1st?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

a + c − 4b − c + a − b a + a + c - c -4b - b

OpenStudy (anonymous):

@surdawi i'm not really sure how to do this....

OpenStudy (anonymous):

a+a=?

OpenStudy (anonymous):

b ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

a + a just like numbers 1 + 1 =2 a + a = 2a c - c ?

OpenStudy (anonymous):

0c ?

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

0*c= 0 correct?

OpenStudy (anonymous):

-4b - b

OpenStudy (anonymous):

-5b ?

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

so we have 2a + 0c -5b which answer?

OpenStudy (anonymous):

2a-5b @surdawi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much!! you're a lot of help!!

OpenStudy (anonymous):

@surdawi